OpenStudy (anonymous):

Find a general solution solution of the homogeneous system [1 2 1; -2 1 -7; 3 4 5][x1 x2 x3] = [0 0 0]

myininaya (myininaya):

this is what i got

OpenStudy (anonymous):

I got the same! so, the solution should look like: |-3*x3 | x= | x3 | | x3 | it's one column matrix where is x3 is any number

OpenStudy (anonymous):

Or you can write it as\[x = \left[ \begin{array}{c} -3\\1 \\1 \end{array} \right]x_3\implies x \in Span \{\left[ \begin{array}{c} -3\\1 \\1 \end{array} \right]\}\]

myininaya (myininaya):

yes i forgot the pretty things that we can write gj inink and polpak